use long division to find the quotient and the remainder when the function f(x) = x³+3x²-4 is divided by (x-1).hence factorize the function completely
1
Expert's answer
2020-11-26T15:27:04-0500

"\begin{matrix}
 & x^2 +4x+4 \ \ \\
 \ \ \ \ \ \ \ \ \ \ \ \ \ x-1 &)\overline{x^3+3x^2-4} \\
& - \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\
& \underline{x^3-x^2} \ \ \ \ \ \ \ \ \\
& \ \ \ \ \ \ 4x^2-4 \\
& \ \ - \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\
& \ \ \ \ \ \ \ \ \underline{4x^2-4x} \\
& \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 4x-4 \\
& \ \ - \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\
& \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \underline{4x-4} \\
& \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 0 \\
\end{matrix}"

Therefore,


"\dfrac{x^3+3x^2-4}{x-1}=x^2+4x+4=(x+2)^2"

"x^3+3x^2-4=(x-1)(x+2)^2"